# MuLan: Multimodal-LLM Agent for Progressive Multi-Object Diffusion

### v1

@@ Appendix A Detailed prompt template of the global planning by the LLM
Appendix B Detailed prompt template of the local planning by the LLM
@@
Appendix C More details on the overlapping processing
@@
$\bm{M}_{n,i}=\begin{cases}\Big{(}\tilde{x}_{n-1}\cdot r_{i}+(\tilde{x}_{n-1}+\tilde{w}_{n-1})\cdot(1-r_{i}),\tilde{y}_{n-1},\tilde{w}_{n-1}\cdot r_{i}+\frac{W-\tilde{x}_{n-1}-\tilde{w}_{n-1}}{\texttt{Num}^{n}},\tilde{h}_{n-1}\Big{)},\text{if }\texttt{opt}^{n}=\texttt{right},\\ \\ \Big{(}\tilde{x}_{n-1},\frac{(\texttt{Num}^{n}-1)\cdot\tilde{y}_{n-1}}{\texttt{Num}^{n}},\tilde{w}_{n-1},\tilde{h}_{n-1}\cdot r_{i}+\frac{\tilde{y}_{n-1}}{\texttt{Num}^{n}}\Big{)},\text{if }\texttt{opt}^{n}=\texttt{top}.\end{cases}$ (8)
@@
Appendix D More details on the evaluation questionnaire
As shown in Section 4, we design a questionnaire to comprehensively evaluate the alignment between the generated image and the text by GPT-4V (OpenAI, 2023) and human, from three aspects - object completeness, correctness of attribute bindings, and correctness of spatial relationships. Specifically, given an image and a text prompt, for object completeness, we will evaluate if the image contains each single object in the prompt. If the object appears in the image, we will then judge if the attribute bindings of the object in the image align with the corresponding attribute bindings in the text prompt, to evaluate the correctness of attribute bindings. We will also ask GPT-4V or human to judge if the spatial relationships are correct and match the text, as the evaluation of the spatial relationships.
Examples of the questionnaire for different images and text prompts are shown in Figure 8.
Figure 8: Illustration of the questionnaire for the evaluation of generated images
Appendix E More qualitative results

### v2

@@ Appendix E Detailed prompt template of the global planning by the LLM
Appendix F Detailed prompt template of the local planning by the LLM
@@
Appendix G Details for the computation of rough masks
When $n=1$, since there is no object generated yet, both the position $\texttt{opt}_{1}$ and $\texttt{Num}_{1}$ are unrestricted and the LLM can be prompted to determine $\texttt{opt}_{1}$ and $\texttt{Num}_{1}$ given sub-prompt $\texttt{p}_{1}$. Since the object order starts from left to right and bottom to top, there will be only two position options $\texttt{opt}_{1}\in\{\texttt{left},\texttt{bottom}\}$ for $\texttt{obj}_{1}$. Once $\texttt{opt}_{1}$ determined, MuLan evenly splits the whole image’s width/height ($W/H$) to $\texttt{Num}_{1}$ parts and assigns the very left (bottom) part to $\texttt{obj}_{1}$, which leads to the following bounding box (an illustration for the computation is shown in Figure 6):
$\bm{M}_{1}=\begin{cases}(0,0,\frac{W}{\texttt{Num}_{1}},H),&\text{if }\texttt{opt}_{1}=\texttt{left},\\ (\frac{(\texttt{Num}_{1}-1)\cdot H}{\texttt{Num}_{1}},0,W,\frac{H}{\texttt{Num}_{1}}),&\text{if }\texttt{opt}_{1}=\texttt{bottom}.\end{cases}$ (5)
[…]
$\bm{M}_{n}=\begin{cases}(\tilde{x}_{n-1}+\tilde{w}_{n-1},0,\frac{W-\tilde{x}_{n-1}+\tilde{w}_{n-1}}{\texttt{Num}_{n}},H),&\text{if }\texttt{opt}_{n}=\texttt{right},\\ \\ (0,\frac{\tilde{y}_{n-1}\cdot(\texttt{Num}_{n}-1)}{\texttt{Num}_{n}},W,\frac{\tilde{y}_{n-1}}{\texttt{Num}_{n}}),&\text{if }\texttt{opt}_{n}=\texttt{top}.\end{cases}$ (6)
Appendix H More details on the overlapping processing
@@
$\bm{M}_{n,i}=\begin{cases}\Big{(}\tilde{x}_{n-1}\cdot r_{i}+(\tilde{x}_{n-1}+\tilde{w}_{n-1})\cdot(1-r_{i}),\tilde{y}_{n-1},\tilde{w}_{n-1}\cdot r_{i}+\frac{W-\tilde{x}_{n-1}-\tilde{w}_{n-1}}{\texttt{Num}_{n}},\tilde{h}_{n-1}\Big{)},\\ \text{if }\texttt{opt}_{n}=\texttt{right},\\ \\ \Big{(}\tilde{x}_{n-1},\frac{(\texttt{Num}_{n}-1)\cdot\tilde{y}_{n-1}}{\texttt{Num}_{n}},\tilde{w}_{n-1},\tilde{h}_{n-1}\cdot r_{i}+\frac{\tilde{y}_{n-1}}{\texttt{Num}_{n}}\Big{)},\\ \text{if }\texttt{opt}_{n}=\texttt{top}.\end{cases}$ (7)
@@
Appendix I More details on the evaluation questionnaire
As shown in Section 4, we design a questionnaire to comprehensively evaluate the alignment between the generated image and the text by GPT-4V [15] and human, from three aspects - object completeness, correctness of attribute bindings, and correctness of spatial relationships. Specifically, given an image and a text prompt, for object completeness, we will evaluate if the image contains each single object in the prompt. If the object appears in the image, we will then judge if the attribute bindings of the object in the image align with the corresponding attribute bindings in the text prompt, to evaluate the correctness of attribute bindings. We will also ask GPT-4V or human to judge if the spatial relationships are correct and match the text, as the evaluation of the spatial relationships.
Examples of the questionnaire for different images and text prompts are shown in Figure LABEL:fig:evaluation.
Appendix J More qualitative results